Consider the following. x = 5 - 6t y = 6 + 5t (a) Sketch the curve presented parametric equations (indicate the orientation c I need the solve answer B
Answer #1

SOL:- x = 5-6€ - y = 6151 it in they and put of t, in terms first writ a 1 = 5-6€ t-s=-6t 672 5- y = 675+ = 65 (5) 34+ (25-54